OpenStudy (anonymous):

the sum of two numbers is 7/8. One addend is 3/4 . What is the other

OpenStudy (nfcfox):

Well let's get common denominators so we can solve. This is like saying 3/4+x= 7/8 and x is what number we are looking for. In order to solve this we need common denominators. 3/4 is also 6/8 as we can multiply the denominator by 2 to get 8 so we also multiply the numerator by 2. The equation is now 6/8+x=7/8. Subtract 6/8 and we get x= 1/8. Our other number is 1/8.

OpenStudy (anonymous):

ok thx